OpenStudy (anonymous):

what is the horizontal asymptote of f(x)= 1/ (3+x^2)

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

x=infinity so 1/infinity approaches 0 answer is 0

OpenStudy (anonymous):

with your eyeballs. degree of denominator is bigger than degree of the numerator

OpenStudy (anonymous):

?

OpenStudy (anonymous):

infinity is not a number. neither is \[\frac{1}{\infty}\]

OpenStudy (anonymous):

limit as the equation approaches infinity is approaching 0

OpenStudy (anonymous):

here is a nice picture. you can see that as x gets bigger \[\frac{1}{x^2+3}\] gets smaller http://www.wolframalpha.com/input/?i=1%2F%28x^2%2B3%29+domain+-10+to+10